# GPU-Accelerated Gravitational Lensing &\&& Dynamical (GLaD) Modeling for Cosmology and Galaxies

### v1

@@ 3.4 Bayesian information criterion (BIC)
In this section, we introduce a BIC method to distinguish the goodness of mass models of lens galaxies with different $\bm{\eta_{\rm LD}}$. The BIC is an approximation to the Bayesian evidence
@@
where $k$ is the number of parameters in the model, $n$ is the number of data points, and $\mathcal{L}$ is the maximum likelihood of the model. The BIC penalizes models with a higher number of parameters, effectively balancing goodness of fit with model simplicity. The likelihood in our case is the product of the lensing modeling $\mathcal{L}(\bm{d}_{\rm L}\leavevmode\nobreak\ |\leavevmode\nobreak\ \bm{\eta}_{\rm LD})$ and dynamical modeling $\mathcal{L}(\bm{d}_{\rm D}\leavevmode\nobreak\ |\leavevmode\nobreak\ \bm{\eta}_{\rm LD})$. The likelihood is easily overwhelmed by the lensing data due to the large amount of pixels on the extended arcs. In this work, we focus on using spatially resolved kinematics data to break the internal MSD and constrain $\lambda_{\rm int}$. The lensing-only modeling cannot constrain $\lambda_{\rm int}$. Thus, we discard $\mathcal{L}(\bm{d}_{\rm L}\leavevmode\nobreak\ |\leavevmode\nobreak\ \bm{\eta}_{\rm LD})$ and only make use of the difference of $\mathcal{L}(\bm{d}_{\rm D}\leavevmode\nobreak\ |\leavevmode\nobreak\ \bm{\eta}_{\rm LD})$ from the joint modeling to weight the posterior distribution.
@@
with ${\rm BIC}_{i}-{\rm BIC}_{\rm min}>=0$. As discussed in Sect. 3.1, the preferred lensing mass parameters vary across different parameter spaces depending on the source resolution. The choice of source pixelization introduces uncertainties in the BIC for a given lens mass parametrization (see Appendix. C). To quantify this uncertainty, we compare the BIC values across different source grids and measure the root-mean-square scatter $\sigma_{\rm BIC}$. Following Birrer et al. (2019) and Yıldırım et al. (2020), we incorporate this uncertainty into the model weighting by convolving $f_{\rm BIC}$ in Eq. 56 with a Gaussian distribution of variance $\sigma_{\rm BIC}^{2}$, thereby obtaining the updated model weights:
@@ 4 Simulated mock datasets
RXJ1131 was discovered by Sluse et al. (2003). The lens galaxy is located at a redshift of $z_{\text{lens}}=0.295$, while the lensed source galaxy is at a redshift of $z_{\text{s}}=0.654$, both confirmed through spectroscopy (e.g., Sluse et al. 2007). The lens is accompanied by a faint satellite galaxy S (see Fig. 3), which JWST NIRSpec has confirmed to be at the same redshift as the lens (see Shajib et al., in prep). Imaging data was collected from the Hubble Space Telescope (HST) Advanced Camera for Surveys (ACS) with an exposure time of 1980 seconds. Time-delay measurements for RXJ1131 were made through a dedicated optical monitoring campaign under the COSMOGRAIL program (e.g., Tewes et al. 2013). These measurements, based on frequent observations (every 3 days) over more than 9 years and involving over 700 epochs using meter-class telescopes and new curve-shifting techniques, reported an approximately $3\%$ precision time delay by Tewes et al. (2013); Liao et al. (2015); Bonvin et al. (2017). Microlensing-induced time-delay shifts, as analyzed by Tie & Kochanek (2018), have been found to be negligible within the context of the extended delay, as discussed by Chen et al. (2018).
@@
To simulate the kinematics map, we follow the approach presented in Yıldırım et al. (2020). We use the best-fit lensing light map for the kinematic mock data and assume a Poisson noise-dominated region. The relative pixel intensities are then converted into a relative 2D signal-to-noise map. We adopt VorBin [^10] https://pypi.org/project/vorbin/ package (Cappellari & Copin 2003) to apply the adaptive spatial binning to the signal-to-noise ratio map, with a target signal-to-noise ratio of 50 per bin. We simulate the data with a high signal-to-noise ratio to ensure that by combining high-quality kinematic data, the internal MSD can be effectively broken. Considering the light contamination from nearby quasar images and the extended host galaxy at the Einstein radius of $\theta_{\rm E}\simeq 1.65\leavevmode\nobreak\ \arcsec$, the simulated binned map covers a small field of view (FoV) ranging from $-1\arcsec$ to $1\arcsec$ relative to the lens centroid (see Fig. 3). For simplicity, we neglect the satellite when mocking up the IFU map as well as during the modeling of the SL and stellar kinematic data. We assume a single Gaussian kinematic $\rm PSF_{\rm kin}$ with a FWHM of $0.14\arcsec$, which corresponds approximately to the PSF measured from JWST NIRSpec data of RXJ1131 (see Shajib et al., in prep). We generate the noiseless kinematic map with JamPy [^11] https://pypi.org/project/jampy/ package based on the mass and light distribution from the best-fit lens model (refer to the best-fit parameters in Tab. 3) and the simulated binned map.
@@
Table 3: Model parameters and prior for joint modeling. The value of $D_{\rm d}$ is determined from $z_{\rm d}=0.295$, assuming $H_{0}=82.5\leavevmode\nobreak\ \rm km\leavevmode\nobreak\ s^{-1}\leavevmode\nobreak\ Mpc^{-1}$ , $\Omega_{\rm m}=0.27$ and $\Omega_{\Lambda}=0.73$. On the contrary, the value of $D_{\rm\Delta t,int}$ must be corrected for external convergence (see Eq. 11) to obtain the true distance. Applying this correction, we find $D_{\Delta\rm t,int,ext}=1980.14$ Mpc, given the assumed cosmology. The position angle $\theta_{\rm PA}$ defines the orientation of both the light and mass profiles of the lens galaxy, measured counterclockwise from the +$x$-axis and is fixed during modeling process. The scale radius $r_{\rm s}$ in the dark matter profile indicates the slope transition of the density profile from $-1$ (inner region) to $-3$ (outer region) which cannot be well constrained both by SL and kinematic data due to its large distance from the galaxy centroid. For this reason, a Gaussian prior is used in the joint modeling (Gavazzi et al. 2007). In each joint modeling, the BH mass is fixed, but we explore different models by probing $M_{\rm BH}$ within the range of [$10^{9},10^{10}]$ ${\rm M_{\odot}}$. Note that $H_{0}$ and $\Omega_{\rm m}$ are not directly modeled in the joint analysis. The prior in the table indicates the range from which the samples are drawn and then evaluated by the posterior of the distances.
Figure 3: Mock data sets of lensing imaging and kinematic data. Note that we observe a faint satellite galaxy above the lens galaxy at the same redshift as the lens galaxy. We neglect the satellite when mocking up the IFU data. The satellite is too small to extract useful kinematic information from the IFU datacube other than the redshift (see Shajib et al. in prep.). More importantly, based on the previous study, the satellite has a negligible effect on mass modeling and cosmological distance inference (Suyu et al. 2014). The mock IFU data with 52 bins is presented in the same reference frame as the simulated HST imaging, with north oriented upwards and east to the left.
@@ 5 Analysis and discussion of the joint modeling results
In this section, we present the results of the joint modeling using mock lensing and kinematic data. In Sect. 5.1, we discuss the fitting results of the joint modeling and demonstrate how it breaks the internal MSD, given ideal kinematic data. In Sect. 5.2, we examine the impact of black hole mass on $H_{0}$ inference, given ideal kinematic data. In Sect. 5.3, we present the joint modeling, given the ideal kinematic excluding the central region to probe if the impact of an unknown BH mass can be mitigated. In Sect. 4, we analyze the effect of systematic errors in the kinematic map on $H_{0}$.
@@ 5.1 Breaking the MSD using joint modeling
We perform joint modeling using ideal kinematic data (see Eq. 60). Based on the velocity dispersion of the lens galaxy in RXJ1131, measured as $\sigma_{\rm disp}=320\pm 20$ km/s in Suyu et al. (2014), and the $M_{\rm BH}-\sigma_{\rm disp}$ relation, we explore the full range of possible BH masses of $[10^{9}\leavevmode\nobreak\ {\rm M_{\odot}},10^{10}\leavevmode\nobreak\ {\rm M_{\odot}}]$ to be conservative.
We adopt the composite mass model $\kappa_{\rm int,comp}$ and perform joint modeling over the same black hole mass range. In joint modeling, we fix the BH mass in a given model setup and increment it in steps of of $10^{9}\leavevmode\nobreak\ {\rm M_{\odot}}$ within the $M_{\rm BH}$ range across multiple model setups. Additionally, we perform joint modeling by replacing the composite mass model with a single EPL mass profile, $\kappa_{\rm int,epl}$. Throughout the modeling process, we do not allow $M_{\rm BH}$ to vary. Each run is performed with a fixed $M_{\rm BH}$ on a source grid ranging from $60\times 60$ to $68\times 68$ pixels, increasing in steps of 2 to account for the degeneracy caused by source-grid resolutions. These source resolutions are sufficient to address parameter degeneracies while achieving a good fit for the extended arcs (see details in Appendix C). From the equal-weighted probability density of $D_{\rm d}$, we observe that $D_{\rm d}$ is broadly distributed across the prior range. Models with the same $M_{\rm BH}$ but different source grid resolutions form tightly clustered distributions of $D_{\rm d}$. When considering different $M_{\rm BH}$ values, the $D_{\rm d}$ distribution accounting for degeneracy from source grid resolution variations, spans a wide range from 650 Mpc to 850 Mpc. In contrast, $D_{\rm\Delta t,int}$ is more tightly clustered around the input value and remains almost unaffected by $M_{\rm BH}$ (see Fig. 4). This behavior is expected since $D_{\rm d}$ is primarily constrained by dynamical modeling, making it more sensitive to $M_{\rm BH}$. Fig. 5 illustrates more clearly the relation between $D_{\rm d}$ and $M_{\rm BH}$ that was shown in Fig. 4.
Figure 4: Top panels: Marginalized posterior density distribution of $D_{\rm d}$, based on joint models using ideal kinematic data. Different colors represent posterior densities corresponding to different black hole masses, while the different colors indicate models with identical mass parameterization but different source grid resolutions. The red color represents EPL mass models, where a small softening scale of $r_{\rm soft}=0.01\arcsec$ mimics the presence of a massive BH, eliminating the need to explicitly include an additional $M_{\rm BH}$. The left panel shows equally weighted $D_{\rm d}$ posterior densities, whereas the right panel presents the combined $D_{\rm d}$ posterior density weighted by BIC (see Sect. 3.4). Bottom panels: Marginalized posterior density distribution of $D_{\rm\Delta t,int}$, based on joint models using ideal kinematic data. The red dashed lines in both panels indicate the mock input values used in the simulated data. The black dashed lines represent the median values in the BIC weighted distribution.
Figure 5: $M_{\rm BH}$ versus $D_{\rm d}$ with $1\sigma$ uncertainties, based on the ideal kinematic data. This provides a clearer visualization of the upper-left panel in Fig. 4. The plot also illustrates how the BH mass affects the inference of $D_{\rm d}$, given that $\beta_{\rm ani}$ is within the prior range of $-0.3$ to $0.3$. The gray shaded region represents models with the BH mass satisfying $f_{\rm BIC}^{*}\geq 0.2$, indicating a significant contribution to $D_{\rm d}$ when combining all models together.
The time-delay distance $D_{\rm\Delta t,int}$ is entirely degenerate with $\lambda_{\rm int}$ over the prior range of $\lambda_{\rm int}$ when considering lensing-only modeling. The kinematic data aid in constraining $\lambda_{\rm int}$ and in identifying the uniquely preferred $\kappa_{\rm int}$ model within the range $\lambda_{\rm int}\in[0.5,1.5]$. Consequently, we can break the internal MSD and firmly constrain $D_{\Delta\rm t,int}$ (see the red box in Fig. 6). We combine all joint modelings across different mass model assumptions, values of $M_{\rm BH}$, and source-grid resolutions, weighting them using $\mathcal{L}(\bm{d}_{\rm D}\leavevmode\nobreak\ |\leavevmode\nobreak\ \bm{\eta}_{\rm LD})$ within the BIC framework (see Sect. 3.4). Models where $M_{\rm BH}$ deviates significantly from the mock input in the simulated kinematic data obtain lower weights. Additionally, the EPL model exhibits a higher scatter in the probability density distribution for $D_{\rm\Delta t,int}$ across the different source resolutions. As a result, $\lambda_{\rm int}$ is not well constrained in this case, since the kinematic data struggle to differentiate the scaled $\kappa_{\rm int,epl}$. However, $\kappa_{\rm int,epl}$ model is disfavored by BIC weighting, as it fails to accurately reproducing the ideal kinematic data, with $\Delta\chi_{\rm dyn}^{2}=8$ compared to the best-fit composite mass model (see Fig. 7). Ultimately, the recovered time-delay distance is $D_{\rm\Delta t,int}=1857_{-78}^{+137}$ Mpc, which deviates from the mock input by $1.87\%$, within the 1-sigma uncertainty range. Similarly, the recovered lens distance $D_{\rm d}=781_{-29}^{+30}$ Mpc shows a deviation of $0.77\%$.
@@
The uncertainty in $D_{\rm\Delta t,int}$ is asymmetric, exhibiting a longer tail on the positive side and a shorter tail on the negative side (see Figs. 4 and 6). This occurs because, as $\lambda_{\rm int}$ approaches the upper bound of 1.5 in the prior, it implies that $\kappa_{\rm gal}$ is being modified by the addition of a negative constant sheet $(1-\lambda_{\rm int})$ on top of $\lambda_{\rm int}\kappa_{\rm gal}$ (see Eq. 6). At regions far from the lensing centroid, $\kappa_{\rm int}$ becomes negative, which is disallowed by dynamical modeling in the framework of JAM.
A perfect mass profile for $\lambda_{\rm int}$ to characterize internal MSD would ideally remain constant up to a distance of $\sim 20\arcsec$ to the lens centroid, where it is largely insensitive to lensing data of RXJ1131 but can still capture changes in the mass density slope induced by internal MSD through kinematic data. Beyond this distance, it should immediately drop to zero. Therefore, if the sheet is perfect, the modeled $\kappa_{\rm int}$ will remain non-negative up to $50\arcsec$ in our setup and will not be rejected by JAM (see Sect. 3.3). To approximate the internal MSD, we use a dPIE profile, which declines rapidly to zero beyond the truncation radius but does not exhibit a strict cut-off. This relatively gradual decline in the range of $20\arcsec$ to $50\arcsec$ results in a region where $\kappa_{\rm int}$ becomes negative, leading to an asymmetric probability distribution for $D_{\rm\Delta t,int}$. Consequently, the lower $1\sigma$ bound of 78 Mpc may be underestimated compared to the true $1\sigma$ interval, assuming $\lambda_{\rm int}$ behaves as an idealized sheet with an abrupt cut-off beyond $\sim 20\arcsec$, as described above.
The probability distribution of $D_{\rm d}$ is also slightly asymmetric, but it is less pronounced than that of $D_{\Delta\rm t,int}$. The asymmetry of $D_{\rm d}$ arises from the influence of $M_{\rm BH}$. As $M_{\rm BH}$ becomes heavier, $D_{\rm d}$ tends to shift towards lower values and extends down to 650 Mpc to accommodate the kinematic data. Since we use $\mathcal{L}(\bm{d}_{\rm D}\leavevmode\nobreak\ |\leavevmode\nobreak\ \bm{\eta}_{\rm LD})$ for BIC weighting in joint models with varying $M_{\rm BH}$, some models with larger $M_{\rm BH}$ can achieve a similarly good fit to the kinematic data as the model corresponding to the true $M_{\rm BH}$ by appropriately rescaling the distances $D_{\rm d}$ and the anisotropy parameter $\beta_{\rm ani}$. These models contribute to the tails of the inferred $D_{\rm d}$ distribution (see Fig. 5 and the upper panels in Fig. 4).

### v2

@@ 3.4 Bayesian information criterion (BIC)
In this section, we introduce a BIC method to distinguish the goodness of mass models of lens galaxies with different $\bm{\eta_{\rm LD}}$. The BIC is an approximation to the Bayesian evidence.
@@
where $k$ is the number of parameters in the model, $n$ is the number of data points, and $\mathcal{L}$ is the maximum likelihood of the model. The BIC penalizes models with a higher number of parameters, effectively balancing goodness of fit with model simplicity. The likelihood in our case is the product of the lensing modeling $\mathcal{L}(\bm{d}_{\rm L}\leavevmode\nobreak\ |\leavevmode\nobreak\ \bm{\eta}_{\rm LD})$ and dynamical modeling $\mathcal{L}(\bm{d}_{\rm D}\leavevmode\nobreak\ |\leavevmode\nobreak\ \bm{\eta}_{\rm LD})$. The likelihood is easily overwhelmed by the lensing data due to the large number of pixels on the extended arcs. In this work, we focus on using spatially resolved kinematics data to break the internal MSD and constrain $\lambda_{\rm int}$. The lensing-only modeling cannot constrain $\lambda_{\rm int}$ or distinguish between a composite and an elliptical power-law model. Therefore, we exclude $\mathcal{L}(\bm{d}_{\rm L}\leavevmode\nobreak\ |\leavevmode\nobreak\ \bm{\eta}_{\rm LD})$ and the lensing data points, relying on $\mathcal{L}(\bm{d}_{\rm D}\leavevmode\nobreak\ |\leavevmode\nobreak\ \bm{\eta}_{\rm LD})$, the number of kinematic data points, and the number of parameters $\bm{\eta}_{\rm LD}$ adopted in the joint modeling to weight the posterior distribution.
@@
with ${\rm BIC}_{i}-{\rm BIC}_{\rm min}>=0$. As discussed in Sect. 3.1, the preferred lensing mass parameters vary across different parameter spaces depending on the source resolution. The choice of source pixelization introduces uncertainties in the BIC for a given lens mass parametrization (see Appendix. D). To quantify this uncertainty, we compare the BIC values across different source grids and measure the root-mean-square scatter $\sigma_{\rm BIC}$. Following Birrer et al. (2019) and Yıldırım et al. (2020), we incorporate this uncertainty into the model weighting by convolving $f_{\rm BIC}$ in Eq. 52 with a Gaussian distribution of variance $\sigma_{\rm BIC}^{2}$, thereby obtaining the updated model weights:
@@ 4 Simulated mock datasets
RXJ1131 was discovered by Sluse et al. (2003). The lens galaxy is located at a redshift of $z_{\text{lens}}=0.295$, while the lensed source galaxy is at a redshift of $z_{\text{s}}=0.654$, both confirmed through spectroscopy (e.g., Sluse et al. 2007). The lens is accompanied by a faint satellite galaxy S (see Fig. 3), which JWST NIRSpec has confirmed to be at the same redshift as the lens (see Shajib et al. (2025)). Imaging data was collected from the Hubble Space Telescope (HST) Advanced Camera for Surveys (ACS) with an exposure time of 1980 seconds. Time-delay measurements for RXJ1131 were made through a dedicated optical monitoring campaign under the COSMOGRAIL program (e.g., Tewes et al. 2013). These measurements, based on frequent observations (every 3 days) over more than 9 years and involving over 700 epochs using meter-class telescopes and new curve-shifting techniques, reported an approximately $3\%$ precision time delay by Tewes et al. (2013); Liao et al. (2015); Bonvin et al. (2017). Microlensing-induced time-delay shifts, as analyzed by Tie & Kochanek (2018), have been found to be negligible within the context of the extended delay, as discussed by Chen et al. (2018).
@@
To simulate the kinematics map, we follow the approach presented in Yıldırım et al. (2020). We use the best-fit lensing light map for the kinematic mock data and assume a Poisson noise-dominated region. The relative pixel intensities are then converted into a relative 2D signal-to-noise map. We adopt VorBin [^9] https://pypi.org/project/vorbin/ package (Cappellari & Copin 2003) to apply the adaptive spatial binning to the signal-to-noise ratio map, with a target signal-to-noise ratio of 50 per bin. We simulate the data with a high signal-to-noise ratio to ensure that by combining high-quality kinematic data, the internal MSD can be effectively broken. Considering the light contamination from nearby quasar images and the extended host galaxy at the Einstein radius of $\theta_{\rm E}\simeq 1.65\leavevmode\nobreak\ \arcsec$, the simulated binned map covers a small field of view (FoV) ranging from $-1\arcsec$ to $1\arcsec$ relative to the lens centroid (see Fig. 3). For simplicity, we neglect the satellite when mocking up the IFU map as well as during the modeling of the SL and stellar kinematic data. We assume a single Gaussian kinematic $\rm PSF_{\rm kin}$ with a FWHM of $0.14\arcsec$, which corresponds approximately to the PSF measured from JWST NIRSpec data of RXJ1131 (see Shajib et al. (2025)). We generate the noiseless kinematic map with JamPy [^10] https://pypi.org/project/jampy/ package based on the mass and light distribution from the best-fit lens model (refer to the best-fit parameters in Tab. 11) and the simulated binned map.
@@
Table 3: Model parameters and prior for joint modeling
@@
 [^11] The value of $D_{\rm d}$ is determined from $z_{\rm d}=0.295$, assuming $H_{0}=82.5\leavevmode\nobreak\ \rm km\leavevmode\nobreak\ s^{-1}\leavevmode\nobreak\ Mpc^{-1}$ , $\Omega_{\rm m}=0.27$ and $\Omega_{\Lambda}=0.73$. On the contrary, the value of $D_{\rm\Delta t,int}$ must be corrected for external convergence (see Eq. 12) to obtain the true distance. Applying this correction, we find $D_{\Delta\rm t,int}=1980.14$ Mpc, given the assumed cosmology. The position angle $\theta_{\rm PA}$ defines the orientation of both the light and mass profiles of the lens galaxy, measured counterclockwise from the +$x$-axis and is fixed during modeling process. The scale radius $r_{\rm s}$ in the dark matter profile indicates the slope transition of the density profile from $-1$ (inner region) to $-3$ (outer region), which cannot be well constrained by either SL or kinematic data due to its large distance from the galaxy centroid. For this reason, a Gaussian prior is used in the joint modeling (Gavazzi et al. 2007). In each joint modeling, the BH mass is fixed, but we explore different models by probing $M_{\rm BH}$ within the range of [$10^{9},10^{10}]$ ${\rm M_{\odot}}$. Note that $H_{0}$ and $\Omega_{\rm m}$ are not directly modeled in the joint analysis. The prior in the table indicates the range from which the samples are drawn and then evaluated by the posterior of the distances.
Figure 3: Mock data sets of lensing imaging and kinematic data. Note that we observe a faint satellite galaxy above the lens galaxy at the same redshift as the lens galaxy. We neglect the satellite when mocking up the IFU data. The satellite is too small to extract useful kinematic information from the IFU datacube other than the redshift (see Shajib et al. (2025)). More importantly, based on the previous study, the satellite has a negligible effect on mass modeling and cosmological distance inference (Suyu et al. 2014). The mock IFU data with 52 bins is presented in the same reference frame as the simulated HST imaging, with north oriented upwards and east to the left.
@@ 5 Analysis and discussion of the joint modeling results
In this section, we present the results of the joint modeling using mock lensing and kinematic data. In Sect. 5.1, we outline the joint modeling setup and describe the modeling procedure. In Sect. 5.2, we discuss the fitting results of the joint modeling and demonstrate how it breaks the internal MSD, given ideal kinematic data. In Sect. 5.3, we analyze the effect of systematic errors in the kinematic map on $H_{0}$, given the $5\%$ biased kinematic datasets. In Sect. 5.4, we examine the impact of $M_{\rm BH}$ on $H_{0}$ inference, given ideal kinematic data. In Sect. 5.5, we present joint modeling based on ideal kinematic data, excluding the central region, to test whether the impact of an unknown black hole mass can be mitigated, thereby reducing potential bias in distance measurements. We also discuss how the $M_{\rm BH}$$\beta_{\rm ani}$ degeneracy has been addressed in the literature and explore the potential for applying these solutions to lensing and dynamical modeling.
5.1 Joint modeling setup and procedure
The mass models adopted in our joint modeling are $\kappa_{\rm int,comp}$ and $\kappa_{\rm int,epl}$ (see Sect. 3.3). Both explicitly represent the mass sheet using a dPIE profile. In the composite model, each component of the galaxy is modeled separately with distinct mass profiles. To account for the central BH, we consider the lens galaxy in RXJ1131, whose velocity dispersion has been measured as $\sigma_{\rm disp}=320\pm 20$ km s-1 by Suyu et al. (2014). Using the $M_{\rm BH}-\sigma_{\rm disp}$ relation, we explore the full range of possible BH masses of $[10^{9}\leavevmode\nobreak\ {\rm M_{\odot}},10^{10}\leavevmode\nobreak\ {\rm M_{\odot}}]$ to be conservative. For each composite mass model setup, we fix the BH mass and increment it in steps of $10^{9}{\rm M_{\odot}}$ across the specified range. In contrast, the EPL mass model treats the galaxy as a single mass component and does not include an additional mass profile for the BH.
For each mass model setup, whether using $\kappa_{\rm int,comp}$ or $\kappa_{\rm int,epl}$, we perform joint modeling across a range of source grids, from $60\times 60$ to $68\times 68$ pixels, increasing in steps of 2. This variation accounts for potential degeneracies introduced by the source-grid resolution. These resolutions are sufficient to mitigate parameter degeneracies while ensuring a good fit to the extended arcs (see Appendix D for details).
The lensing constraints in the joint modeling are consistently provided by the same simulated lensing data. For the kinematic input, we consider two sets of simulated data: one ideal (see Eq. 56) and one with a $5\%$ bias (see Eq. 57). For each case, we perform 55 joint modeling runs (1 EPL model plus 10 composite models for the 10 fixed BH masses, and each of the 11 models has 5 source-grid resolutions). The final posterior distributions of the lensing and dynamical parameters are then obtained by combining the results of these 55 models using BIC weighting (see Sect. 3.4), for the ideal and $5\%$ biased kinematic datasets, respectively.
@@ 5.2 Breaking the MSD using joint modeling
In this section, we illustrate how joint modeling resolves the internal MSD given an ideal kinematic dataset. The time-delay distance $D_{\rm\Delta t,int}$ is entirely degenerate with $\lambda_{\rm int}$ over the prior range of $\lambda_{\rm int}$ when considering lensing-only modeling. The kinematic data aid in constraining $\lambda_{\rm int}$ and in identifying the uniquely preferred $\kappa_{\rm int}$ model within the range $\lambda_{\rm int}\in[0.5,1.5]$. Consequently, we can break the internal MSD and firmly constrain $D_{\Delta\rm t,int}$ (see the red box in Fig. 4).
We present both equal-weighted and BIC-weighted histograms of $D_{\rm d}$ and $D_{\rm\Delta t,\mathrm{int}}$ in Fig. 5. The BIC does not clearly distinguish between composite mass models with BH masses in the range $M_{\mathrm{BH}}=2\times 10^{9},M_{\odot}$ to $5\times 10^{9},M_{\odot}$, as the differences in BIC values $\Delta\mathrm{BIC}$ are comparable to their associated uncertainties $\sigma_{\mathrm{BIC}}$ (see Appendix F). Outside this mass range, as well as for the EPL model, the BIC is more effective at discriminating between models. The limited effectiveness of BIC weighting is consistent with expectations, as $M_{\rm BH}$ cannot be well constrained by dynamical modeling for galaxies beyond the local universe. The EPL model, in particular, exhibits higher scatter in the probability density distribution of $D_{\rm\Delta t,\mathrm{int}}$ across different source resolutions. This increased scatter is likely due to the fact that the mock kinematic data were generated using a composite mass model. Since the EPL model assumes a single power-law mass profile, it fails to fully capture the complexity of the true lens mass distribution, leading to less reliable constraints on $D_{\rm\Delta t,\mathrm{int}}$. Despite this internal scatter, the EPL model is statistically disfavored under BIC weighting. It consistently underperforms in fitting the mock kinematic data. We obtain a minimal discrepancy of $\Delta\chi_{\rm dyn}^{2}=8$ of the EPL model across different source grid resolutions relative to the best-fit composite mass model (see Fig. 6).
We combine all 55 models, obtaining the recovered time-delay distance $D_{\rm\Delta t,int}=1857_{-78}^{+137}$ Mpc, which deviates from the mock input by $1.87\%$, within the 1-sigma uncertainty range. Similarly, the recovered lens distance $D_{\rm d}=781_{-29}^{+30}$ Mpc shows a deviation of $0.77\%$.
@@
The uncertainty in $D_{\rm\Delta t,int}$ is asymmetric, exhibiting a longer tail on the positive side and a shorter tail on the negative side (see Figs. 4 and 5). This occurs because, as $\lambda_{\rm int}$ approaches the upper bound of 1.5 in the prior, it implies that $\kappa_{\rm gal}$ is being modified by the addition of a negative constant sheet $(1-\lambda_{\rm int})$ on top of $\lambda_{\rm int}\kappa_{\rm gal}$ (see Eq. 6). In regions far from the lensing centroid, $\kappa_{\rm int}$ becomes negative, which is not allowed under the dynamical modeling framework of JAM. As a result, the probability distribution of $D_{\rm\Delta t,int}$ becomes asymmetric (see Sect. 3.3). The lower $1\sigma$ bound of 78 Mpc may be underestimated relative to the true $1\sigma$ interval, assuming that $\lambda_{\rm int}$ follows an “ideal” mass sheet profile with a sharp cutoff beyond $\sim 20\arcsec$, which mimics the MST and also allows higher values of $\lambda_{\rm int}$ without leading to negative $\kappa_{\rm int}$. The probability distribution of $D_{\rm d}$ is also slightly asymmetric, but it is less pronounced than that of $D_{\Delta\rm t,int}$. The asymmetry in $D_{\rm d}$ arises from the influence of $M_{\rm BH}$, which is discussed in more detail in Sect. 5.4.
@@
Figure 5: Top panels: Marginalized posterior density distribution of $D_{\rm d}$, based on joint models using ideal kinematic data. Different colors represent posterior densities corresponding to different BH masses, while each color represents five separate posterior distributions that are tightly clustered together, corresponding to models with identical mass parameterizations but different source-grid resolutions. The red color represents EPL mass models, where a small softening scale of $r_{\rm soft}=0.01\arcsec$ mimics the presence of a massive BH, eliminating the need to explicitly include an additional $M_{\rm BH}$. The left panel shows equally weighted $D_{\rm d}$ posterior densities, whereas the right panel presents the combined $D_{\rm d}$ posterior density weighted by BIC (see Sect. 3.4). Bottom panels: Marginalized posterior density distribution of $D_{\rm\Delta t,int}$, based on joint models using ideal kinematic data. The red dashed lines in both panels indicate the mock input values used in the simulated data. The black dashed lines represent the median values in the BIC weighted distribution.
@@
Figure 8: Marginalized posterior density distribution of $\beta_{\rm ani}$, based on joint models using ideal kinematic data. Different colors represent posterior densities corresponding to different BH masses, while the same color indicates models with identical mass parameterization but different source grid resolutions. We observe that as $M_{\rm BH}$ increases, the inferred $\beta_{\rm ani}$ decreases and vice versa. The inferred $\beta_{\rm ani}$ distributions given different $M_{\rm BH}$ spread over the prior range [--0.3,0.3], but different $M_{\rm BH}$ yield different goodness of fit to the kinematic data (as shown in Fig. 5).
Figure 9: $M_{\rm BH}$ versus $D_{\rm d}$ with $1\sigma$ uncertainties, based on the ideal kinematic data. This provides a clearer visualization of the upper-left panel in Fig. 5. The plot also illustrates how the BH mass affects the inference of $D_{\rm d}$, given that $\beta_{\rm ani}$ is within the prior range of $-0.3$ to $0.3$. The grey shaded region represents models with the BH mass satisfying $f_{\rm BIC}^{*}\geq 0.2$, indicating a significant contribution to $D_{\rm d}$ when combining all models together.
The 5% biased kinematic data also helps break the internal MSD, yielding consistent results for $\lambda=0.97_{-0.07}^{+0.04}$ and $D_{\rm\Delta t,int}=1863_{-80}^{+144}$ Mpc, which agree with values inferred from the joint modeling using ideal kinematic data. This demonstrates that the overall systematic bias does not affect the constraints on $D_{\rm\Delta t,int}$ and $\lambda_{\rm int}$ (see Fig. 4). This is because $\lambda_{\rm int}$ is constrained by the 2D kinematic map, where the shape of the $v_{\rm rms}$ profile breaks the internal MSD and constrains $D_{\rm\Delta t,int}$. The 5% bias does not alter the shape of the $v_{\rm rms}$ profile, which is why neither $D_{\rm\Delta t,int}$ nor $\lambda_{\rm int}$ is biased. The orbital anisotropy parameter, $\beta_{\rm ani}$, is primarily constrained by the shape of the kinematic map; therefore, its inference remains robust. We obtain $\beta_{\rm ani}=0.20^{+0.08}_{-0.13}$, which is consistent with the value inferred from the ideal kinematic data.
The systematic bias primarily impacts $D_{\rm d}$ because it changes the amplitude of the $v_{\rm rms}$ overall. Given the relation $v_{\rm rms}^{\rm pre}\sim\frac{1}{\sqrt{D_{\rm d}}}$, a $5\%$ bias in $v_{\rm rms}^{\rm pre}$ results in an expected $\sim 9\%$ bias in $D_{\rm d}$. We obtain $D_{\rm d}=706_{-25}^{+20}$ Mpc, which is $9\%$ lower than the mock input value of $D_{\rm d}=775$ Mpc, as expected. If the combined kinematics are obtained from a single aperture rather than an IFU, the impact on distances will not be cleanly isolated to $D_{\rm d}$ alone, as the single aperture lacks information on the shape of $v_{\rm rms}$. We anticipate a more severe effect on both $D_{\rm d}$ and $D_{\rm\Delta t,int}$.